Eliminate the arbitrary constant a from 2z=(ax+y)^2
1
Expert's answer
2019-08-22T11:48:10-0400

We will use the following standard notation to denote the partial derivatives:


"{\eth z \over \eth x}=p, {\eth z \over \eth y}=q"

Given "2z=(ax+y)^2"

Differentiating the equation partially with respect to x and y respectively we get


"2{\eth z \over \eth x}=2a(ax+y)"

"2{\eth z \over \eth y}=2(ax+y)"

Then


"p=a(ax+y), q=ax+y"

We have


"px+qy-q^2=ax(ax+y)+y(ax+y)-(ax+y)^2="

"=(ax+y)^2-(ax+y)^2=0"

We show that "2z=(ax+y)^2", where "a" is an arbitrary constant, is a complete integral of


"px+qy-q^2=0"

The partial differential equation is


"2z=q^2"